OpenStudy (anonymous):

I got this answer wrong in my homework; "A body moving in the positive x direction passes the origin at time t = 0. Between t = 0 and t = 1 second, the body has a constant speed of 15 meters per second. At t = 1 second, the body is given a constant acceleration of 5 meters per second squared in the negative x direction. The position of the body at t = 10 seconds is..." Apparently the correct answer is -53m... I keep getting -67m. Why is it -53m? Can someone show me please?

OpenStudy (anonymous):

 let x2= x0 and use this equation : X2-X0=V0t - .5a(t^2) X2-15=15*9-.5*5*81 ,,,,, let t= 9 >>>>> t=10-1 X2=-53